A typical raindrop is much more massive than a mos
Answer #1

(a)

given that

U = 8.2 m/s

we know

M*U = (M + m)V

where M = 40 *m

so

V = M*U / (M+m)

V = 40*m*8.2 /41m

V = 8m/s

(b)

given that

dT = 8ms= 8*10^(-3) s

we know that

F = m*dV/dT

F = m*8/8*10^(-3)

F = 1000m

and

from above relationship

a = 1000 m/s^2

a = 1000/g = 1000/ 9.8 = 102.04 g's